OpenStudy (anonymous):

f(x)=x^2-3 and g(x)=2x-1, how do I find the value of g(f(3))? Step by step please!

OpenStudy (anonymous):

first find \[f(3)\] and \[f(3)=3^2-3=9-3=6\]

OpenStudy (anonymous):

then find \[g(6)=2\times 6-1=12-1=11\]

OpenStudy (anonymous):

and therefore \[g(f(3))=g(6)=11\]

OpenStudy (anonymous):

the correct answer was 2x^2-7, I need the value please.

OpenStudy (anonymous):

\[g(f(3))\] is not a function, it is a number. perhaps you meant \[g(f(x))\]

OpenStudy (anonymous):

oh yes, it was g(f(x)) i'm sorry, my mistake can you help me with that one?

OpenStudy (anonymous):

yes we work as follows \[g(f(x))=g(x^2-3)\] i replaced \[f(x)\] by \[x^2-3\] now replace the "x" in \[g(x)\] by \[x^2-3\] and get \[g(x^2-3)=2\times (x^2-3)-1\] and now some algebra to finish up \[2\times (x^2-3)-1=2x^2-6-1=2x^2-7\]

OpenStudy (anonymous):

Oh I see now, thank you!